### v2

# Robust targeted exploration for systems with non-stochastic disturbances

Abstract
In this paper, we introduce a novel targeted exploration strategy designed specifically for uncertain linear time-invariant systems with energy-bounded disturbances, i.e., without making any assumptions on the distribution of the disturbances. We use classical results characterizing the set of non-falsified parameters consistent with energy-bounded disturbances. We derive a semidefinite program which computes an exploration strategy that guarantees a desired accuracy of the parameter estimate. This design is based on sufficient conditions on the spectral content of the exploration data that robustly account for initial parametric uncertainty. Finally, we highlight the applicability of the exploration strategy through a numerical example involving a nonlinear system.
keywords:
Experiment design, Robust estimation, Optimization under uncertainties, Uncertainty Quantification, Data-driven control
thanks: Frank Allgöwer is thankful that his work was funded by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2075 - 390740016 and under grant 468094890. Frank Allgöwer acknowledges the support by the Stuttgart Center for Simulation Science (SimTech). Janani Venkatasubramanian thanks the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting her. Johannes Köhler was supported by the Swiss National Science Foundation under NCCR Automation (grant agreement 51NF40 180545).
1 Introduction
Designing reliable controllers for unknown dynamical systems requires accurate knowledge of the model parameters, which can be obtained from data [gevers2005identification]. The accuracy of the parameters significantly depends on the quality of the data used for system identification. Informative data can be strategically obtained from an experiment through the process of targeted exploration or optimal experiment design [pronzato2008optimal, gevers1986optimal]. Specifically, targeted exploration inputs are tailored to reduce model uncertainty, thereby ensuring the attainment of a desired accuracy in the identified model [jansson2005input, bombois2021robust], or the feasibility of robust control design  [barenthin2008identification, umenberger2019robust, ferizbegovic2019learning, venkatasubramanian2020robust, venkatasubramanian2023sequential]. In this paper, we develop a targeted exploration strategy for uncertain linear systems subject to energy-bounded disturbances.
Targeted exploration is typically studied in a stochastic setup, for example, with independent and identically distributed (i.i.d.) disturbances with zero mean. In this case, one can construct a data-based confidence ellipsoid for the parameters [ljung1999system] that can be approximately predicted and optimized. This classical asymptotic result has been utilized to design targeted exploration strategies for dual control methods [barenthin2008identification, umenberger2019robust, iannelli2019structured, venkatasubramanian2020robust, venkatasubramanian2023sequential]. In [dean2017sample], tight confidence regions for the parameters are constructed, assuming the data is independent, and hence it is not applicable to correlated time-series data from a single trajectory. Nevertheless, this bound has been utilized in the design of targeted exploration for dual control in [ferizbegovic2019learning].
A common feature among all of the discussed targeted exploration approaches is that they consider linear systems subject to i.i.d. zero mean stochastic disturbances. However, real-world systems often exhibit nonlinear behaviour or unmodeled dynamics, i.e., errors in the assumed model structure, which introduce additional deterministic model mismatch and cannot be explained by independent stochastic noise [sarker2023accurate]. Instead, such disturbances can be modeled as bounded disturbances, assuming that the disturbances belong to a known bounded set, e.g., point-wise bounds or energy bounds. To quantify models under such disturbances, set-membership estimation methods have been developed, and various identification results in this direction can be found in [fogel1979system, milanese1989estimation, sarker2023accurate, bisoffi2021trade, shakouri2024identification]. Although these methods have recently gained popularity for data-driven robust control design [van2020noisy, berberich2022combining], a principled method for optimal experiment design in the presence of bounded disturbances is still lacking.
In this paper, we design a targeted exploration strategy that ensures a desired error bound on the estimated parameters by utilizing a data-dependent uncertainty bound based on energy-bounded disturbances [fogel1979system]. In contrast to existing optimal experiment design methods, which often rely on stochastic disturbance assumptions [umenberger2019robust, ferizbegovic2019learning, venkatasubramanian2023sequential], our approach provides robustness against non-stochastic, adversarial disturbances and is applicable to systems with bounded nonlinearities. We consider multi-sine exploration inputs of specific frequencies and optimized amplitudes to explicitly shape and reduce uncertainty in a targeted manner. As one of our main contributions, we derive a sufficient condition on the spectral content of the exploration data that asymptotically guarantees a desired error-bound on the parameters estimated through exploration. We utilize the sufficient condition on the spectral content to derive LMIs for exploration which ensure the desired error bound on the parameters. This approach gives rise to a targeted exploration design with minimal input energy based on a semidefinite program (SDP). While existing work has also derived SDPs for targeted exploration design [venkatasubramanian2023sequential], the non-stochastic identification results considered in the proposed work yield structurally different requirements on excitation, which require new worst-case robust guarantees rather than high-probability guarantees. To this end, we provide the setting and define the exploration goal in Section 2, and the preliminaries regarding data-dependent uncertainty bounds in Section 3. In Section 4, we derive the exploration strategy by leveraging a sufficient condition on the time-series exploration data, as proposed in our previous work [venkatasubramanian2023towards]. Furthermore, we robustly account for parametric uncertainty and error due to disturbances with suitable bounds in Section 4, unlike [venkatasubramanian2023towards]. Finally, in Section 5, we provide a numerical example to highlight the applicability of the proposed exploration strategy to nonlinear systems.
2 Problem statement
Notation: The transpose of a matrix $A\in\mathbb{R}^{n\times m}$ is denoted by $A^{\top}$. The conjugate transpose of a matrix $A\in\mathbb{C}^{n\times m}$ is denoted by $A^{\mathsf{H}}$. The positive definiteness and positive semi-definiteness of a matrix $A\in\mathbb{C}^{n\times n}$ is denoted by $A=A^{\mathsf{H}}\succ 0$ and $A=A^{\mathsf{H}}\succeq 0$, respectively. The operator $\mathrm{vec}(A)$ stacks the columns of $A$ to form a vector. The operator $\textnormal{diag}(A_{1},\dots,A_{n})$ creates a block diagonal matrix by aligning the matrices $A_{1},\dots,A_{n}$ along the diagonal starting with $A_{1}$ in the upper left corner. The Kronecker product is denoted by $\otimes$. The Euclidean norm and the weighted Euclidean norm for a vector $x\in\mathbb{R}^{n}$ and a matrix $P\succ 0$ are denoted by $\|x\|=\sqrt{x^{\top}x}$ and $\|x\|_{P}=\sqrt{x^{\top}Px}$, respectively. The largest singular value of a matrix $A\in\mathbb{C}^{m\times n}$ is denoted by $\|A\|$. Furthermore, given a matrix $M\succeq 0$, $\|A\|_{M}=\|M^{1/2}A\|$ where $M^{1/2}$ is the symmetric square root matrix of $M$. The identity matrix of size $n$ is denoted by $I_{n}$. A vector of ones of size $n$ is denoted by $\mathbf{1}_{n}\in\mathbb{R}^{n\times 1}$.
Amplitude of a spectral line [venkatasubramanian2023sequential, Def. 2]: Given a sequence $\{\phi_{k}\}_{k=0}^{T-1}$, the amplitude of the spectral line of the sequence at a frequency $\omega\in\Omega_{T}:=\{0,1/T,\dots,(T-1)/T\}$ is given by
$\bar{\phi}(\omega):=\frac{1}{T}\sum_{k=0}^{T-1}\phi_{k}e^{-j2\pi\omega k}.$ (1)
2.1 Setting
Consider a discrete-time, linear time-invariant system of the form
$x_{k+1}=A_{\mathrm{tr}}x_{k}+B_{\mathrm{tr}}u_{k}+w_{k},$ (2)
where $x_{k}\in\mathbb{R}^{n_{\mathrm{x}}}$ is the state, $u_{k}\in\mathbb{R}^{n_{\mathrm{u}}}$ is the control input, and $w_{k}\in\mathbb{R}^{n_{\mathrm{x}}}$ is the disturbance. In our setting, the true system parameters $A_{\mathrm{tr}}$, $B_{\mathrm{tr}}$ are uncertain. Hence, it is necessary to collect informative data from an optimal experiment for a fixed $T\in\mathbb{N}$ time steps to enhance the accuracy of the parameters. It is assumed that the state can be measured, the initial state is at the origin, i.e., $x_{0}=0$, and the disturbances are energy-bounded.
Assumption 1.
The disturbances $w$ are energy-bounded, i.e., there exists a known constant $\gamma_{\mathrm{w}}>0$ such that
$\sum_{k=0}^{T-1}\|w_{k}\|^{2}\leq\gamma_{\mathrm{w}}.$ (3)
This assumption on the disturbances allows the system to exhibit nonlinear behaviour, as long as the nonlinearities can be bounded in energy (as discussed in Section 5). Importantly, this assumption represents the main difference to existing experiment design methods which are typically restricted to i.i.d. Gaussian disturbances, e.g., [umenberger2019robust, ferizbegovic2019learning, venkatasubramanian2023sequential].
Exploration goal: Since the true system parameters $\theta_{\mathrm{tr}}=\text{vec}([A_{\mathrm{tr}}$, $B_{\mathrm{tr}}])\in\mathbb{R}^{n_{\theta}}$, with $n_{\theta}=n_{\mathrm{x}}(n_{\mathrm{x}}+n_{\mathrm{u}})$, are not precisely known, exploratory inputs should be designed to excite the system to gather informative data. Specifically, our objective is to design inputs that excite the system in a manner as to obtain an estimate $\hat{\theta}_{T}=\text{vec}([\hat{A}_{T},\hat{B}_{T}])$ that satisfies
$(\theta_{\mathrm{tr}}-\hat{\theta}_{T})^{\top}(D_{\mathrm{des}}\otimes I_{n_{\mathrm{x}}})(\theta_{\mathrm{tr}}-\hat{\theta}_{T})\leq 1,$ (4)
where $D_{\mathrm{des}}\succ 0$ is a user-defined matrix characterizing closeness of $\hat{\theta}_{T}$ to $\theta_{\mathrm{tr}}$. We assume that we have some prior knowledge about the system dynamics.
Assumption 2.
The unknown parameters $\theta_{\mathrm{tr}}=\mathrm{vec}([A_{\mathrm{tr}},B_{\mathrm{tr}}])$ lie in a known set $\mathbf{\Theta}_{0}$, i.e., $\theta_{\mathrm{tr}}\in\mathbf{\Theta}_{0}$, where
$\mathbf{\Theta}_{0}:=\left\{\theta:(\hat{\theta}_{0}-\theta)^{\top}(D_{0}\otimes I_{n_{\mathrm{x}}})(\hat{\theta}_{0}-\theta)\leq 1\right\},$ (5)
with an estimate $\hat{\theta}_{0}=\mathrm{vec}([\hat{A}_{0},\hat{B}_{0}])$ for some $D_{0}\succ 0$.
Remark 3.
An initial estimate $\hat{\theta}_{0}$ and set $\mathbf{\Theta}_{0}$ as in Assumption 2 result naturally from a finite-horizon experiment under Assumption 1 (cf. Lemma 6).
2.2 Exploration strategy
The exploration input sequence takes the form
$u_{k}=\sum_{i=1}^{L}\bar{u}(\omega_{i})\cos(2\pi\omega_{i}k),\quad\,k=0,\dots,T-1$ (6)
where $T$ is the exploration time and $\bar{u}(\omega_{i})\in\mathbb{R}^{n_{\mathrm{u}}}$ are the amplitudes of the sinusoidal inputs at $L\in\mathbb{N}$ distinct selected frequencies $\omega_{i}\in\Omega_{T}$ with $n_{\mathrm{x}}+n_{\mathrm{u}}\leq L\leq T$. In practice, frequencies may be selected based on prior information [mehra1974optimal]. Since the input signal is deterministic and sinusoidal, the amplitude of the spectral line of the sequence $\{u_{k}\}_{k=0}^{T-1}$ at frequency $\omega_{i}$ is $\bar{u}(\omega_{i})$. Denote $U_{\mathrm{e}}=\mathrm{diag}(\bar{u}(\omega_{1}),\dots,\bar{u}(\omega_{L}))\in\mathbb{R}^{Ln_{\mathrm{u}}\times L}$. The exploration input is computed such that it excites the system sufficiently with minimal input energy, based on the initial parameter estimates. Bounding the input energy by a constant $\gamma^{2}_{\mathrm{e}}$ can be equivalently written as $\sum_{i=1}^{L}\|\bar{u}(\omega_{i})\|^{2}=\mathbf{1}_{L}^{\top}U_{\mathrm{e}}^{\top}U_{\mathrm{e}}\mathbf{1}_{L}\preceq\gamma_{\mathrm{e}}^{2}$ where $\mathbf{1}_{L}\in\mathbb{R}^{L\times 1}$ is a vector of ones, and the bound $\gamma_{\mathrm{e}}\geq 0$ is desired to be small. Using the Schur complement, this criterion is equivalent to
$S_{\textnormal{energy-bound}}(\gamma_{\mathrm{e}},U_{\mathrm{e}})\coloneqq\begin{bmatrix}\gamma_{\mathrm{e}}&\mathbf{1}_{L}^{\top}U_{\mathrm{e}}^{\top}\\ U_{\mathrm{e}}\mathbf{1}_{L}&\gamma_{\mathrm{e}}I\end{bmatrix}\succeq 0.$ (7)
In order to design the exploration inputs, we make the following assumption regarding the system dynamics.
Assumption 4.
The system matrix $A_{\mathrm{tr}}$ is Schur stable.
Remark 5.
We require $A_{\mathrm{tr}}$ to be Schur stable since we consider only open-loop inputs in our exploration strategy (6). Assumption 4 could be relaxed if an exploration input of the form in (6) with an additional linear feedback, i.e., $v_{k}=u_{k}+Kx_{k}$, is utilized which ensures robust stability for all $\theta\in\mathbf{\Theta}_{0}$ (5).
In order to achieve the exploration goal, the amplitudes of the sinusoidal exploration inputs need to be optimized such that by applying the exploration inputs, the obtained estimate satisfies the desired uncertainty bound (4).
3 Preliminaries on data-driven uncertainty quantification
In this section, we discuss a data-dependent uncertainty bound on the parameter estimates in the presence of energy-bounded noise [fogel1979system]. Given observed data $\mathcal{D}_{T+1}=\{x_{k},u_{k}\}_{k=0}^{T}$ of length $T+1$, the objective is to quantify the uncertainty associated with the unknown parameters $\theta_{\mathrm{tr}}$. Henceforth, we denote $\phi_{k}=[x_{k}^{\top}\,u_{k}^{\top}]^{\top}\in\mathbb{R}^{n_{\phi}}$ where $n_{\phi}=n_{\mathrm{x}}+n_{\mathrm{u}}$. The system (2) can be re-written in terms of parameter $\theta_{\mathrm{tr}}=\textnormal{vec}([A_{\mathrm{tr}},B_{\mathrm{tr}}])$ as
$x_{k+1}=(\phi_{k}^{\top}\otimes I_{n_{\mathrm{x}}})\theta_{\mathrm{tr}}+w_{k}.$ (8)
In order to simplify the exposition, we denote
$\Phi=[\phi_{0},\dots,\phi_{T-1}]\in\mathbb{R}^{n_{\phi}\times T}$ (9)
and
$X^{\top}=[x_{1}^{\top},\dots,x_{T}^{\top}]\in\mathbb{R}^{1\times Tn_{\mathrm{x}}}.$ (10)
We obtain the following expressions for the mean $\hat{\theta}_{T}~=~\textnormal{vec}([\hat{A}_{T},\hat{B}_{T}])$ and covariance $P$ of the parameters from the standard least squares formulation [ljung1999system, Section 1.3]:
$\hat{\theta}_{T}=P\sum_{k=0}^{T-1}(\phi_{k}^{\top}\otimes I_{n_{\mathrm{x}}})^{\top}x_{k+1}=P(\Phi\otimes I_{n_{\mathrm{x}}})X$ (11)
and
$P=\left(\sum_{k=0}^{T-1}\phi_{k}\phi_{k}^{\top}\right)^{-1}\otimes I_{n_{\mathrm{x}}}=(\Phi\Phi^{\top})^{-1}\otimes I_{n_{\mathrm{x}}}.$ (12)
The non-falsified region for the uncertain parameters $\theta$ is provided in the following lemma.
Lemma 6.
[fogel1979system] Let Assumption 1 hold. Given data set $\mathcal{D}_{T+1}$, the set of non-falsified parameters $\theta$ is given by
$\mathbf{\Theta}_{T}:=\left\{\theta:(\theta-\hat{\theta}_{T})^{\top}{P^{-1}}(\theta-\hat{\theta}_{T})\leq G\right\}$ (13)
where
$G=\gamma_{\mathrm{w}}+\|\hat{\theta}_{T}\|_{P^{-1}}^{2}-X^{\top}X.$ (14)
{pf}
The energy constraint on the process noise in (2) yields the following non-falsified set:
$\mathbf{\Theta}_{T}$ $=\left\{\theta:\sum_{k=0}^{T-1}\|x_{k+1}-(\phi_{k}^{\top}\otimes I_{n_{\mathrm{x}}})\theta\|^{2}\leq\gamma_{\mathrm{w}}\right\}$ (15)
which can be equivalently written as
$\theta^{\top}\left((\Phi\Phi^{\top})\otimes I_{n_{\mathrm{x}}}\right)\theta-2\left(X^{\top}(\Phi^{\top}\otimes I_{n_{\mathrm{x}}})\right)\theta$
$\leq$ $\;\gamma_{\mathrm{w}}-X^{\top}X.$ (16)
By adding $\|\hat{\theta}_{T}\|_{P^{-1}}^{2}$ to both sides of (3), and using (11) and (12) to complete the squares, we get
$\|\theta-\hat{\theta}_{T}\|_{P^{-1}}^{2}\leq\gamma_{\mathrm{w}}+\|\hat{\theta}_{T}\|_{P^{-1}}^{2}-X^{\top}X=:G,$ (17)
[…]
Given Assumption 1, the non-falsified set $\bm{\Theta}_{T}$ provides an exact characterization of the set of parameters explaining the data. Similar non-falsified sets for matrices are considered in [van2020noisy, berberich2022combining]. The ellipsoid (13), derived from energy-bounded constraints, is characterized by a vector $\hat{\theta}_{T}$ and a matrix $P$, which correspond to the mean and covariance of the least squares estimator for linear systems with Gaussian disturbances [umenberger2019robust, Prop. 2.1]. However, unlike the case of zero mean i.i.d. Gaussian disturbances, in the case of energy-bounded disturbances the scaling $G$ (14) of the bounding ellipsoid is also data-dependent. This dependence has significant consequences on optimal experiment design, as will be discussed in the next section. Furthermore, for $\gamma_{\mathrm{w}}\propto T$, $G$ scales at most linearly with $T$ as $T\to\infty$ since $G\leq\gamma_{\mathrm{w}}$ [fogel1979system, Lemma 4]. In contrast, the scaling of the confidence ellipsoid in the Gaussian case does not depend on $T$ [ljung1999system, umenberger2019robust]. Since $P^{-1}$ increases linearly with $T$, the size of the confidence ellipsoid in the Gaussian case reduces with $T$ [ljung1999system, umenberger2019robust]. However, in the considered case of energy-bounded disturbances, the size of the non-falsified set $\mathbf{\Theta}_{T}$, in general, does not decrease as $T\to\infty$.
4 Targeted Exploration
In this section, we propose a targeted exploration strategy based on the data-dependent uncertainty bound provided in Lemma 6. The exploration strategy builds upon sufficient conditions on the exploration data outlined in Section 4.1. In particular, we derive bounds on the exploration data using the spectral information of the exploration inputs in Section 4.2. Since these bounds depend on the uncertain model parameters, bounds on the effect of model uncertainty are derived in Section 4.3. We utilize these bounds to derive sufficient conditions on the spectral information of the exploration inputs. However, since the derived sufficient conditions are non-convex in the decision variables, a convex relaxation procedure is carried out in Section 4.5. Finally, in Section 4.6, the exploration problem is reduced to a set of LMIs that provide us exploration inputs that ensure the exploration goal.
4.1 Sufficient conditions for exploration
Given the form of the exploration inputs in (6), the exploration goal (4), and the data-dependent uncertainty bound in Lemma 6, in what follows, we provide conditions that the exploration data have to satisfy to achieve the exploration goal. Denote the Cholesky decomposition of $D_{\mathrm{des}}$ as $D_{\mathrm{des}}=D_{\mathrm{des}}^{\frac{1}{2}\top}D_{\mathrm{des}}^{\frac{1}{2}}$. The following proposition presents a sufficient condition to ensure that the exploration goal is achieved.
Theorem 7.
[venkatasubramanian2023towards, Theorem 4] Suppose $\Phi$ and $X$ satisfy
$\begin{bmatrix}\Phi\Phi^{\top}-\gamma_{\mathrm{w}}D_{\mathrm{des}}&0\\ 0&0\end{bmatrix}$
$+\underbrace{\begin{bmatrix}D_{\mathrm{des}}^{\frac{1}{2}\top}(X^{\top}\otimes I_{n_{\phi}})\\ (\Phi\otimes I_{n_{\mathrm{x}}})\otimes I_{n_{\phi}}\end{bmatrix}}_{=:Z}\begin{bmatrix}D_{\mathrm{des}}^{\frac{1}{2}\top}(X^{\top}\otimes I_{n_{\phi}})\\ (\Phi\otimes I_{n_{\mathrm{x}}})\otimes I_{n_{\phi}}\end{bmatrix}^{\top}$ $\succeq 0.$ (18)
Then, the estimate $\hat{\theta}_{T}$ computed as in (11) satisfies the exploration goal (4).
{pf}
The bound in (17) can be re-written as
$(\theta-\hat{\theta}_{T})^{\top}((\Phi\Phi^{\top})\otimes I_{n_{\mathrm{x}}})(\theta-\hat{\theta}_{T})$
[…]
By applying the Schur complement twice to (4.1), we get
$(\theta-\hat{\theta}_{T})(\theta-\hat{\theta}_{T})^{\top}$
$\preceq$ $(\gamma_{\mathrm{w}}-X^{\top}X+X^{\top}((\Phi^{\top}(\Phi\Phi^{\top})^{-1}\Phi)\otimes I_{n_{\mathrm{x}}})X)$
$\cdot(\Phi\Phi^{\top})^{-1}\otimes I_{n_{\mathrm{x}}}.$ (20)
Inequality (7) can be written as
[…]
By applying the Schur complement to (21), we get
$\Phi\Phi^{\top}-D_{\mathrm{des}}^{\frac{1}{2}\top}\left((\gamma_{\mathrm{w}}-X^{\top}X)\otimes I_{n_{\phi}}\right)D_{\mathrm{des}}^{\frac{1}{2}}$ (22)
$-D_{\mathrm{des}}^{\frac{1}{2}\top}\left((X^{\top}((\Phi^{\top}(\Phi\Phi^{\top})^{-1}\Phi)\otimes I_{n_{\mathrm{x}}})X)\otimes I_{n_{\phi}}\right)D_{\mathrm{des}}^{\frac{1}{2}}$ $\succeq 0.$
Since $\Phi\Phi^{\top}\succeq 0\iff(\Phi\Phi^{\top})\otimes I_{n_{\mathrm{x}}}\succeq 0$, (22) can be written as
$(\Phi\Phi^{\top})\otimes I_{n_{\mathrm{x}}}\succeq$ $\big(X^{\top}((\Phi^{\top}(\Phi\Phi^{\top})^{-1}\Phi)\otimes I_{n_{\mathrm{x}}})X$
$+\gamma_{\mathrm{w}}-X^{\top}X\big)(D_{\mathrm{des}}\otimes I_{n_{\mathrm{x}}}).$ (23)
Furthermore, by inserting (4.1) in (4.1), we get
$(\theta-\hat{\theta}_{T})(\theta-\hat{\theta}_{T})^{\top}\preceq(D_{\mathrm{des}}\otimes I_{n_{\mathrm{x}}})^{-1}.$ (24)
[…]
In order to compute the amplitudes of the exploration inputs $U_{\mathrm{e}}$, a few challenges need to be addressed. Note that Inequality (7) depends on $X$ and $\Phi$ quadratically, which further depend on the amplitudes of the exploration inputs $U_{\mathrm{e}}$ (6), as well as the disturbance $w$. Furthermore, since the true dynamics $A_{\mathrm{tr}},\,B_{\mathrm{tr}}$ are uncertain, the linear mapping from the input sequence to the state sequence is not known.
4.2 Bounds based on the theory of spectral lines
In what follows, we address the aforementioned issues by determining sufficient conditions for targeted exploration in terms of the spectral content of $X$ and $\Phi$ based on the theory of spectral lines. Given $u_{k}$ as in (6), $x_{k}$ has $L$ spectral lines from 0 to T1T-1 at distinct frequencies $\omega_{i}\in\Omega_{T}$, $i=1,\dots,L$ with amplitudes [sarker2020parameter, Lemma 1]:
$\bar{x}(\omega_{i})=$ $\underbrace{(e^{j2\pi\omega_{i}}I-A_{\mathrm{tr}})^{-1}B_{\mathrm{tr}}}_{=:V_{\mathrm{x},i}}\bar{u}(\omega_{i})$
$+\underbrace{(e^{j2\pi\omega_{i}}I-A_{\mathrm{tr}})^{-1}}_{=:Y_{\mathrm{x},i}}\bar{w}(\omega_{i})+\bar{x}_{\mathrm{err}}(\omega_{i}).$ (25)
The transient error in the amplitude of a spectral line $\bar{x}_{\mathrm{err}}(\omega_{i})$ decays uniformly (with rate $\frac{1}{\sqrt{T}}$) to 0 as $T\to\infty$ (cf. Assumption 4, [ljung1999system, Theorem 2.1]). To simplify the exposition, we will assume that the transient error can be neglected.
Assumption 8.
The transient error satisfies $\bar{x}_{err}(\omega_{i})=0$ for all $\omega_{i}\in\Omega_{T}$
[…]
$\bar{X}=[\bar{x}(\omega_{1})^{\top},\dots,\bar{x}(\omega_{L})^{\top}]^{\top}\in\mathbb{C}^{n_{\mathrm{x}}L}$ (26)
which satisfies
$\bar{X}=\underbrace{V_{\mathrm{x},\mathrm{tr}}U_{\mathrm{e}}\mathbf{1}_{L}}_{=:\bar{X}_{\mathrm{u}}}+\underbrace{Y_{\mathrm{x},\mathrm{tr}}W\mathbf{1}_{L}}_{=:\bar{X}_{\mathrm{w}}},$ (27)
with
$V_{\mathrm{x},\mathrm{tr}}$ $:=\textnormal{diag}(V_{\mathrm{x},1},\cdots,V_{\mathrm{x},L})\in\mathbb{C}^{n_{\mathrm{x}}L\times n_{\mathrm{u}}L},$
$Y_{\mathrm{x},\mathrm{tr}}$ $:=\textnormal{diag}(Y_{\mathrm{x},1},\cdots,Y_{\mathrm{x},L})\in\mathbb{C}^{n_{\mathrm{x}}L\times n_{\mathrm{x}}L},$
$W$ $:=\mathrm{diag}(\bar{w}(\omega_{1}),\dots,\bar{w}(\omega_{L}))\in\mathbb{C}^{n_{\mathrm{x}}L\times L}.$ (28)
Furthermore, $\phi_{k}$ has $L$ spectral lines from 0 to T1T-1 at distinct frequencies $\omega_{i}\in\Omega_{T}$, $i=1,\dots,L$ with amplitudes
$\bar{\phi}(\omega_{i}):=\underbrace{\begin{bmatrix}V_{\mathrm{x},i}\\ I_{n_{\mathrm{u}}}\end{bmatrix}}_{=:V_{\phi,i}}\bar{u}(\omega_{i})+\underbrace{\begin{bmatrix}Y_{\mathrm{x},i}\\ 0\end{bmatrix}}_{=:Y_{\phi,i}}\bar{w}(\omega_{i}).$ (29)
We compactly define
$\bar{\Phi}=[\bar{\phi}(\omega_{1}),\dots,\bar{\phi}(\omega_{L})]\in\mathbb{C}^{n_{\phi}\times L},$ (30)
which satisfies
$\bar{\Phi}=\underbrace{V_{\mathrm{\phi,tr}}U_{\mathrm{e}}}_{=:\bar{\Phi}_{\mathrm{u}}}+\underbrace{Y_{\mathrm{\phi,tr}}W}_{=:\bar{\Phi}_{\mathrm{w}}},$ (31)
with
$V_{\phi,\mathrm{tr}}$ $:=[V_{\phi,1},\cdots,V_{\phi,L}]\in\mathbb{C}^{n_{\mathrm{\phi}}\times n_{\mathrm{u}}L},$
$Y_{\mathrm{\phi,tr}}$ $:=[Y_{\phi,i},\cdots,Y_{\phi,L}]\in\mathbb{C}^{n_{\mathrm{\phi}}\times n_{\mathrm{x}}L}.$ (32)
The spectral content of $Z$ (7) is denoted by $\bar{Z}\in\mathbb{C}^{(n_{\phi}+n_{\mathrm{x}}n_{\phi}^{2})\times(n_{\mathrm{x}}n_{\phi}L)}$:
$\bar{Z}=$ $\begin{bmatrix}D_{\mathrm{des}}^{\frac{1}{2}\top}\left((\bar{X}_{\mathrm{u}}+\bar{X}_{\mathrm{w}})^{\mathsf{H}}\otimes I_{n_{\phi}}\right)\\ \left((\bar{\Phi}_{\mathrm{u}}+\bar{\Phi}_{\mathrm{w}})\otimes I_{n_{\mathrm{x}}}\right)\otimes I_{n_{\phi}}\end{bmatrix}$
$=$ $\underbrace{\begin{bmatrix}D_{\mathrm{des}}^{\frac{1}{2}\top}\left(\left(\mathbf{1}_{L}^{\top}U_{\mathrm{e}}^{\top}V_{\mathrm{x,tr}}^{\mathsf{H}}\right)\otimes I_{n_{\phi}}\right)\\ \left(V_{\mathrm{\phi,tr}}U_{\mathrm{e}}\right)\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}}_{=:\bar{Z}_{\mathrm{u}}}$
$+\underbrace{\begin{bmatrix}D_{\mathrm{des}}^{\frac{1}{2}\top}\left(\bar{X}_{\mathrm{w}}^{\mathsf{H}}\otimes I_{n_{\phi}}\right)\\ \bar{\Phi}_{\mathrm{w}}\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}}_{=:\bar{Z}_{\mathrm{w}}}.$ (33)
The following lemma provides lower bounds on $\Phi\Phi^{\top}$ and $ZZ^{\top}$ (7) using the spectral content of the signals $x_{k}$ and $\phi_{k}$.
Lemma 9.
Let Assumptions 4 and 8 hold. For any $\epsilon\in(0,1)$, $\phi_{k}$ and $Z$ satisfy
$\Phi\Phi^{\top}\succeq T\left((1-\epsilon)\bar{\Phi}_{\mathrm{u}}\bar{\Phi}_{\mathrm{u}}^{\mathsf{H}}-\left(\frac{1-\epsilon}{\epsilon}\right)\bar{\Phi}_{\mathrm{w}}\bar{\Phi}_{\mathrm{w}}^{\mathsf{H}}\right),$ (34)
and
$ZZ^{\top}\succeq T\left((1-\epsilon)\bar{Z}_{\mathrm{u}}\bar{Z}_{\mathrm{u}}^{\mathsf{H}}-\left(\frac{1-\epsilon}{\epsilon}\right)\bar{Z}_{\mathrm{w}}\bar{Z}_{\mathrm{w}}^{\mathsf{H}}\right),$ (35)
respectively.
The proof of Lemma 9 is provided in Appendix A. The matrices $\bar{\Phi}_{\mathrm{u}}$, $\bar{\Phi}_{\mathrm{w}}$, $\bar{Z}_{\mathrm{u}}$ and $\bar{Z}_{\mathrm{u}}$ in Lemma 9 depend on the transfer matrices $V_{\mathrm{x,tr}}$, $V_{\mathrm{\phi,tr}}$, $Y_{\mathrm{x,tr}}$, and $Y_{\mathrm{\phi,tr}}$. These transfer matrices are dependent on the true dynamics $A_{\mathrm{tr}}$, $B_{\mathrm{tr}}$, and hence, uncertain. Therefore, in what follows, suitable bounds are derived.
4.3 Bounds on transfer matrices
Denote
$\tilde{V}_{\phi}=V_{\mathrm{\phi,tr}}-\hat{V}_{\phi},\;\tilde{V}_{\mathrm{x}}=V_{\mathrm{\mathrm{x},tr}}-\hat{V}_{\mathrm{x}},$ (36)
where the estimates
$\hat{V}_{\phi}$ $=[\hat{V}_{\phi,1},\cdots,\hat{V}_{\phi,L}]\in\mathbb{C}^{n_{\mathrm{\phi}}\times Ln_{\mathrm{u}}},$
$\hat{V}_{\mathrm{x}}$ $=\text{diag}\left(\hat{V}_{\mathrm{x},1},\cdots,\hat{V}_{\mathrm{x},L}\right)\in\mathbb{C}^{n_{\mathrm{\mathrm{x}}}L\times n_{\mathrm{u}}L}$ (37)
are computed using the initial estimates $\hat{\theta}_{0}=\mathrm{vec}([\hat{A}_{0},\hat{B}_{0}])$ (cf. Assumption 2). We can compute matrices $\tilde{\Gamma}_{\phi}$, $\tilde{\Gamma}_{\mathrm{x}}$, $\Gamma_{\mathrm{\phi}}$ and $\Gamma_{\mathrm{\mathrm{x}}}$ such that
$\tilde{V}_{\phi}\tilde{V}_{\phi}^{\mathsf{H}}\preceq\tilde{\Gamma}_{\phi},$ $\;\tilde{V}_{\mathrm{x}}\tilde{V}_{\mathrm{x}}^{\mathsf{H}}\preceq\tilde{\Gamma}_{\mathrm{x}},$ (38)
$Y_{\mathrm{\phi,tr}}Y_{\mathrm{\phi,tr}}^{\mathsf{H}}\preceq\Gamma_{\mathrm{\phi}},$ $\;Y_{\mathrm{\mathrm{x},tr}}Y_{\mathrm{\mathrm{x},tr}}^{\mathsf{H}}\preceq\Gamma_{\mathrm{x}}$ (39)
using $\theta_{\mathrm{tr}}\in\mathbf{\Theta}_{0}$ (cf. Assumption 2). Conditions (38) and (39) are LMIs, and hence matrices $\tilde{\Gamma}_{\mathrm{x}},\tilde{\Gamma}_{\phi}\succ 0$ and $\Gamma_{\phi},\Gamma_{\mathrm{\mathrm{x}}}\succ 0$ may be computed using robust control methods as shown in [venkatasubramanian2023sequential, Appendices A-B], or scenario methods as shown in [venkatasubramanian2023sequential, Appendix C]. Utilizing (39), we derive bounds on $\bar{\Phi}_{\mathrm{w}}$, $\bar{X}_{\mathrm{w}}$ and $\bar{Z}_{\mathrm{w}}$ in the following lemma.
Lemma 10.
Let Assumptions 1, 2, 4 and 8 hold. Given the bounds on $Y_{\phi,\mathrm{tr}}$ and $Y_{\mathrm{x,tr}}$ (39), we have
$\bar{\Phi}_{\mathrm{w}}\bar{\Phi}_{\mathrm{w}}^{\mathsf{H}}\preceq\bar{W}_{\phi}:=\frac{\gamma_{\mathrm{w}}}{T}\Gamma_{\phi}$ (40)
and
$\bar{Z}_{\mathrm{w}}\bar{Z}_{\mathrm{w}}^{\mathsf{H}}$ $\preceq\bar{W}_{Z}:=\left(\frac{\gamma_{\mathrm{w}}}{T}\|\Gamma_{\mathrm{x}}\|\|D_{\mathrm{des}}\|+\frac{\gamma_{\mathrm{w}}}{T}\|\Gamma_{\phi}\|\right)I_{(n_{\phi}+n_{\mathrm{x}}n_{\phi}^{2})}.$ (41)
The proof of Lemma 10 is provided in Appendix B. In what follows, given the bound on transfer matrices, we derive spectral lines-based sufficient conditions for targeted exploration.
4.4 Sufficient conditions for targeted exploration based on the theory of spectral lines
The following proposition provides a condition in terms of the spectral content of $\phi$, which, if satisfied, ensures that the exploration goal (4) is achieved.
Proposition 11.
Let Assumptions 1 and 2 hold. Suppose the matrices $\bar{\Phi}_{\mathrm{u}}$ and $\bar{Z}_{\mathrm{u}}$ satisfy
$\begin{bmatrix}T\left((1-\epsilon)\bar{\Phi}_{\mathrm{u}}\bar{\Phi}_{\mathrm{u}}^{\mathsf{H}}-\left(\frac{1-\epsilon}{\epsilon}\right)\bar{W}_{\phi}\right)-\gamma_{\mathrm{w}}D_{\mathrm{des}}&0\\ 0&0\end{bmatrix}$
$+T\left((1-\epsilon)\bar{Z}_{\mathrm{u}}\bar{Z}_{\mathrm{u}}^{\mathsf{H}}-\left(\frac{1-\epsilon}{\epsilon}\right)\bar{W}_{Z}\right)$ $\succeq 0.$ (42)
Then, the estimate $\hat{\theta}_{T}$ computed as in (11) satisfies the exploration goal (4).
{pf}
Starting from Inequality (11), we have
$0\preceq$ $\begin{bmatrix}T\left((1-\epsilon)\bar{\Phi}_{\mathrm{u}}\bar{\Phi}_{\mathrm{u}}^{\mathsf{H}}-\left(\tfrac{1-\epsilon}{\epsilon}\right)\bar{W}_{\phi}\right)-\gamma_{\mathrm{w}}D_{\mathrm{des}}&0\\ 0&0\end{bmatrix}$
$+T\left((1-\epsilon)\bar{Z}_{\mathrm{u}}\bar{Z}_{\mathrm{u}}^{\mathsf{H}}-\left(\tfrac{1-\epsilon}{\epsilon}\right)\bar{W}_{Z}\right)$
[…]
$+T\left((1-\epsilon)\bar{Z}_{\mathrm{u}}\bar{Z}_{\mathrm{u}}^{\mathsf{H}}-\left(\tfrac{1-\epsilon}{\epsilon}\right)\bar{Z}_{\mathrm{w}}\bar{Z}_{\mathrm{w}}^{\mathsf{H}}\right)$
[…]
Note that $\bar{\Phi}_{\mathrm{u}}$ and $\bar{Z}_{\mathrm{u}}$ depend linearly on the decision variable $U_{\mathrm{e}}$. Determining a lower bound based on the Inequality (11) results in non-convex constraints in $U_{\mathrm{e}}$. To overcome this problem, we utilize a convex relaxation procedure.
4.5 Convex relaxation
The following lemma is utilized to make Inequality (11) linear in the decision variable $U_{\mathrm{e}}$.
Lemma 12.
For any matrices $M\in\mathbb{C}^{n\times m}$ and $N\in\mathbb{C}^{n\times m}$, we have
$MM^{\mathsf{H}}\succeq MN^{\mathsf{H}}+NM^{\mathsf{H}}-NN^{\mathsf{H}}.$ (44)
{pf}
[…]
The following proposition provides a sufficient condition linear in $\bar{Z}_{\mathrm{u}}$ which, ensures the exploration goal (4).
Proposition 13.
Let Assumptions 1 and 2 hold. Suppose the matrices $\bar{\Phi}_{\mathrm{u}}$, $\bar{Z}_{\mathrm{u}}$ and $\hat{Z}\in\mathbb{C}^{(n_{\phi}+n_{\mathrm{x}}n_{\phi}^{2})\times Ln_{\mathrm{x}}n_{\phi}}$ satisfy
$\begin{bmatrix}\left((1-\epsilon)\bar{\Phi}_{\mathrm{u}}\bar{\Phi}_{\mathrm{u}}^{\mathsf{H}}-\left(\frac{1-\epsilon}{\epsilon}\right)\bar{W}_{\phi}\right)-\frac{\gamma_{\mathrm{w}}}{T}D_{\mathrm{des}}&0\\ 0&0\end{bmatrix}$
$+\left((1-\epsilon)\left(\bar{Z}_{\mathrm{u}}\hat{Z}^{\mathsf{H}}+\hat{Z}\bar{Z}_{\mathrm{u}}^{\mathsf{H}}-\hat{Z}\hat{Z}^{\mathsf{H}}\right)-\left(\tfrac{1-\epsilon}{\epsilon}\right)\bar{W}_{Z}\right)$ $\succeq 0,$ (45)
then an estimate $\hat{\theta}_{T}$ computed as in (11) satisfies the exploration goal (4).
{pf}
From Lemma 12, we have
$\bar{Z}_{\mathrm{u}}\bar{Z}_{\mathrm{u}}^{\mathsf{H}}\succeq\bar{Z}_{\mathrm{u}}\hat{Z}^{\mathsf{H}}+\hat{Z}\bar{Z}_{\mathrm{u}}^{\mathsf{H}}-\hat{Z}\hat{Z}^{\mathsf{H}}.$ (46)
[…]
4.6 Exploration SDP
In this section, we provide a sufficient condition that ensure the exploration goal (4) using Proposition 13. In (4.2), $\bar{Z}_{\mathrm{u}}$ can be be written as $\bar{Z}_{\mathrm{u}}=\bar{Z}_{\mathrm{u,1}}+\bar{Z}_{\mathrm{u,2}}$, where
$\bar{Z}_{\mathrm{u,1}}$ $=\begin{bmatrix}D_{\mathrm{des}}^{\frac{1}{2}\top}\left(\mathbf{1}_{L}^{\top}U_{\mathrm{e}}^{\top}\otimes I_{n_{\phi}}\right)\\ 0\end{bmatrix}\left(V_{\mathrm{x,tr}}^{\mathsf{H}}\otimes I_{n_{\phi}}\right),$
$\bar{Z}_{\mathrm{u,2}}$ $=\begin{bmatrix}0\\ \left(V_{\mathrm{\phi,tr}}\right)\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}\left(U_{\mathrm{e}}\otimes I_{n_{\mathrm{x}}n_{\phi}}\right).$ (47)
In order to robustly account for uncertainties in Inequality (13), we split Inequality (13) into three inequalities that are handled separately. The following inequalities are equivalent to Inequality (13) if $\bar{D}_{1}+\bar{D}_{2}+\bar{D}_{3}\succeq 0$:
$(1-\epsilon)\left(\bar{Z}_{\mathrm{u,1}}\hat{Z}^{\mathsf{H}}+\hat{Z}\bar{Z}_{\mathrm{u,1}}^{\mathsf{H}}\right)-\bar{D}_{1}$ $\succeq 0,$ (48a)
$(1-\epsilon)\left(\bar{Z}_{\mathrm{u,2}}\hat{Z}^{\mathsf{H}}+\hat{Z}\bar{Z}_{\mathrm{u,2}}^{\mathsf{H}}-\hat{Z}\hat{Z}^{\mathsf{H}}\right)-\left(\tfrac{1-\epsilon}{\epsilon}\right)\bar{W}_{Z}-\bar{D}_{2}$ $\succeq 0,$ (48b)
$\begin{bmatrix}\left((1-\epsilon)\bar{\Phi}_{\mathrm{u}}\bar{\Phi}_{\mathrm{u}}^{\mathsf{H}}-\left(\frac{1-\epsilon}{\epsilon}\right)\bar{W}_{\phi}\right)-\frac{\gamma_{\mathrm{w}}}{T}D_{\mathrm{des}}&0\\ 0&0\end{bmatrix}-\bar{D}_{3}$ $\succeq 0.$ (48c)
$S_{\textnormal{exp-1}}(\epsilon,\tau_{1},U_{\mathrm{e}},\hat{Z},\hat{V}_{\mathrm{x}},\tilde{\Gamma}_{\mathrm{x}},D_{\mathrm{des}},\bar{D}_{1}):=$ (49a)
$\begin{bmatrix}0&\begin{bmatrix}(1-\epsilon)D_{\mathrm{des}}^{\frac{1}{2}\top}\left(\mathbf{1}_{L}^{\top}U_{\mathrm{e}}^{\top}\otimes I_{n_{\phi}}\right)\\ 0\end{bmatrix}^{\mathsf{H}}\\ \begin{bmatrix}(1-\epsilon)D_{\mathrm{des}}^{\frac{1}{2}\top}\left(\mathbf{1}_{L}^{\top}U_{\mathrm{e}}^{\top}\otimes I_{n_{\phi}}\right)\\ 0\end{bmatrix}&-\bar{D}_{1}\end{bmatrix}-\tau_{1}\begin{bmatrix}-I&(\hat{V}_{\mathrm{x}}^{\mathsf{H}}\otimes I_{n_{\phi}})\hat{Z}^{\mathsf{H}}\\ \hat{Z}(\hat{V}_{\mathrm{x}}\otimes I_{n_{\phi}})&\hat{Z}((\tilde{\Gamma}_{\mathrm{x}}-\hat{V}_{\mathrm{x}}\hat{V}_{\mathrm{x}}^{\mathsf{H}})\otimes I_{n_{\phi}})\hat{Z}^{\mathsf{H}}\end{bmatrix}\succeq 0$
$S_{\textnormal{exp-2}}(\epsilon,\tau_{2},U_{\mathrm{e}},\hat{Z},\bar{W}_{Z},\hat{V}_{\phi},\tilde{\Gamma}_{\phi},D_{\mathrm{des}},\bar{D}_{2}):=$ (49b)
$\begin{bmatrix}0&(1-\epsilon)(U_{\mathrm{e}}\otimes I_{n_{\mathrm{x}}n_{\phi}})\hat{Z}^{\mathsf{H}}\\ (1-\epsilon)\hat{Z}(U_{\mathrm{e}}^{\top}\otimes I_{n_{\mathrm{x}}n_{\phi}})&-(1-\epsilon)\hat{Z}\hat{Z}^{\mathsf{H}}-\left(\tfrac{1-\epsilon}{\epsilon}\right)\bar{W}_{Z}-\bar{D}_{2}\end{bmatrix}-\tau_{2}\begin{bmatrix}-I&\begin{bmatrix}0\\ \hat{V}_{\mathrm{\phi}}\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}^{\mathsf{H}}\\ \begin{bmatrix}0\\ \hat{V}_{\mathrm{\phi}}\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}&\begin{bmatrix}0&0\\ 0&(\tilde{\Gamma}_{\phi}-\hat{V}_{\mathrm{\phi}}\hat{V}_{\mathrm{\phi}}^{\mathsf{H}})\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}\end{bmatrix}\succeq 0$
$S_{\textnormal{exp-3}}(\epsilon,\tau_{3},U_{\mathrm{e}},\hat{U},\bar{W}_{\phi},\hat{V}_{\phi},\tilde{\Gamma}_{\phi},\gamma_{\mathrm{w}},D_{\mathrm{des}},\bar{D}_{3}):=$ (49c)
$\begin{bmatrix}(1-\epsilon)(U_{\mathrm{e}}\hat{U}^{\top}+\hat{U}U_{\mathrm{e}}^{\top}-\hat{U}\hat{U}^{\top})&0\\ 0&\begin{bmatrix}-\left(\tfrac{1-\epsilon}{\epsilon}\right)\bar{W}_{\phi}-\frac{\gamma_{\mathrm{w}}}{T}D_{\mathrm{des}}&0\\ 0&0\end{bmatrix}-\bar{D}_{3}\end{bmatrix}-\tau_{3}\begin{bmatrix}-I&\begin{bmatrix}\hat{V}_{\mathrm{\phi}}\\ 0\end{bmatrix}^{\mathsf{H}}\\ \begin{bmatrix}\hat{V}_{\mathrm{\phi}}\\ 0\end{bmatrix}&\begin{bmatrix}\tilde{\Gamma}_{\phi}-\hat{V}_{\mathrm{\phi}}\hat{V}_{\mathrm{\phi}}^{\mathsf{H}}&0\\ 0&0\end{bmatrix}\end{bmatrix}\succeq 0$
The following theorem provides a sufficient condition linear in $U_{\mathrm{e}}$ which ensures the exploration goal (4).
Theorem 14.
Let Assumptions 1, 2, 4 and 8 hold. Suppose there exist matrices $U_{\mathrm{e}}$, $\bar{D}_{1}$, $\bar{D}_{2}$ and $\bar{D}_{3}$, and scalars $\tau_{1}\geq 0$, $\tau_{2}\geq 0$, and $\tau_{3}\geq 0$ such that
$S_{\textnormal{exp-1}}(\epsilon,\tau_{1},U_{\mathrm{e}},\hat{Z},\hat{V}_{\mathrm{x}},\tilde{\Gamma}_{\mathrm{x}},D_{\mathrm{des}},\bar{D}_{1})\succeq 0,$
$S_{\textnormal{exp-2}}(\epsilon,\tau_{2},U_{\mathrm{e}},\hat{Z},\bar{W}_{Z},\hat{V}_{\phi},\tilde{\Gamma}_{\phi},D_{\mathrm{des}},\bar{D}_{2})\succeq 0,$
$S_{\textnormal{exp-3}}(\epsilon,\tau_{3},U_{\mathrm{e}},\hat{U},\bar{W}_{\phi},\hat{V}_{\phi},\tilde{\Gamma}_{\phi},\gamma_{\mathrm{w}},D_{\mathrm{des}},\bar{D}_{3})\succeq 0,$
$\bar{D}_{1}+\bar{D}_{2}+\bar{D}_{3}\succeq 0,$ (50)
where $S_{\textnormal{exp-1}}$, $S_{\textnormal{exp-2}}$, and $S_{\textnormal{exp-3}}$ are defined in (49a), (49b) and (49c), respectively. Then, an estimate $\hat{\theta}_{T}$ computed as in (11) upon the application of the input (6) satisfies the exploration goal (4).
The proof of Theorem 14 is provided in Appendix C. The key idea of the proof is the application of robust control tools, in particular, the matrix S-lemma [boyd2004convex, vanwaarde2022noisy], to account for parametric uncertainty. In particular, feasibility of LMIs (49b) and (49c) requires $(\tilde{\Gamma}_{\phi}-\hat{V}_{\mathrm{\phi}}\hat{V}_{\mathrm{\phi}}^{\mathsf{H}})\prec 0$, which holds if the initial uncertainty is sufficiently small. Consequently, we can pose the exploration problem of designing exploration inputs that excite the system with minimal energy to obtain an estimate $\hat{\theta}_{T}$ that satisfies (4) using the following SDP:
$\underset{U_{\mathrm{e}},\gamma_{\mathrm{e}},\atop{\tau_{1}\geq 0,\tau_{2}\geq 0,\atop\tau_{3}\geq 0}}{\inf}$ $\;\gamma_{\mathrm{e}}$
s.t. $\;S_{\textnormal{energy-bound}}(\gamma_{\mathrm{e}},U_{\mathrm{e}})\succeq 0$
$\;S_{\textnormal{exp-1}}(\epsilon,\tau_{1},U_{\mathrm{e}},\hat{Z},\hat{V}_{\mathrm{x}},\tilde{\Gamma}_{\mathrm{x}},D_{\mathrm{des}},\bar{D}_{1})\succeq 0$
$\;S_{\textnormal{exp-2}}(\epsilon,\tau_{2},U_{\mathrm{e}},\hat{Z},\bar{W}_{Z},\hat{V}_{\phi},\tilde{\Gamma}_{\phi},D_{\mathrm{des}},\bar{D}_{2})\succeq 0$
$\;S_{\textnormal{exp-3}}(\epsilon,\tau_{3},U_{\mathrm{e}},\hat{U},\bar{W}_{\phi},\hat{V}_{\phi},\tilde{\Gamma}_{\phi},\gamma_{\mathrm{w}},D_{\mathrm{des}},\bar{D}_{3})\succeq 0$
$\;\bar{D}_{1}+\bar{D}_{2}+\bar{D}_{3}\succeq 0.$ (51)
A solution of (4.6) gives us the parameters required for the implementation of the exploration input, i.e., $U_{\mathrm{e}}=\mathrm{diag}(\bar{u}(\omega_{1}),\dots,\bar{u}(\omega_{L}))$, which guarantees the desired uncertainty bound $D_{\mathrm{des}}$ (4). The suboptimality introduced by the convex relaxation procedure can be reduced by iterating Problem (4.6) multiple times until $\gamma_{\mathrm{e}}$ does not change by re-computing $\hat{U}$ and $\hat{Z}$ for the next iteration as
$\hat{U}$ $=U_{\mathrm{e}}^{*},$
$\hat{Z}$ $=\begin{bmatrix}D_{\mathrm{des}}^{\frac{1}{2}\top}\left(\left(\mathbf{1}_{L}^{\top}U_{\mathrm{e}}^{*\top}\hat{V}_{\mathrm{x}}^{\mathsf{H}}\right)\otimes I_{n_{\phi}}\right)\\ \left(\hat{V}_{\mathrm{\phi}}U_{\mathrm{e}}^{*}\right)\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}$ (52)
wherein $U_{\mathrm{e}}^{*}$ is the solution from the previous iteration.
Remark 15.
(Efficient implementation and conservatism) If $D_{\mathrm{des}}$ is a scaled identity matrix, i.e., $D_{\mathrm{des}}=cI_{n_{\phi}}$ with $c>0$, the Kronecker product with $I_{n_{\phi}}$ in LMIs (49a) and (49b) can be factored out. Hence, the dimensions of LMIs (49a) and (49b) can be reduced by a factor of $n_{\phi}$, i.e., to $(Ln_{\mathrm{u}}+1+n_{\mathrm{x}}n_{\phi})$, and to $(Ln_{\mathrm{u}}n_{\mathrm{x}}+1+n_{\mathrm{x}}n_{\phi})$, respectively. The resulting LMIs are equivalent to the LMIs (49a) and (49b) before factoring. Furthermore, LMIs (49a)-(49c) are derived by invoking the matrix S-lemma [van2020noisy]. Enforcing (48a)-(48c) through LMIs (49a)-(49c) does not introduce any conservatism because the satisfaction of LMIs (49a)-(49c) is equivalent to the satisfaction of (48a)-(48c) under conditions (36) and (38).
The overall targeted exploration strategy is summarized in Algorithm 1. The resulting strategy optimally excites the system with exploratory inputs (6) in order to determine model parameters up to a user defined closeness $D_{\mathrm{des}}$.
Algorithm 1 Targeted exploration
1:Specify exploration length $T$, frequencies $\omega_{i},\,i=1,...,L$, energy bound $\gamma_{\mathrm{w}}$, initial estimates $\hat{A}_{0},\,\hat{B}_{0}$, desired accuracy of parameters $D_{\mathrm{des}}$.
2:Compute $\hat{V}_{\phi}$ and $\hat{V}_{\mathrm{x}}$ (4.3) using the initial estimates.
3:Compute bounds $\tilde{\Gamma}_{\phi}$, $\tilde{\Gamma}_{\mathrm{x}}$ (38), and $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$ (39) via methods described in [venkatasubramanian2023sequential, Appendices A-C].
4:Compute matrices $\bar{W}_{\phi}$ (40) and $\bar{W}_{Z}$ (41).
5:Select initial candidates $\hat{Z}$ and $\hat{U}$ (4.6).
6:Set tolerance $\mathrm{tol}>0$.
7:while $\lvert\frac{\gamma_{\mathrm{e}}-\gamma_{\mathrm{e}}^{*}}{\gamma_{\mathrm{e}}}\rvert\geq\mathrm{tol}$ do
8:  Solve the optimization problem (4.6).
9:  Update $\hat{Z}$ and $\hat{U}$ (4.6).
10:end while
11:Apply the exploration input (6) for k=0,,T1k=0,...,T-1.
12:Compute parameter estimate $\hat{\theta}_{T}$ (11); compute parameter set $\bm{\Theta}_{T}$ (13).
4.7 Discussion
In what follows, we examine the key features of the proposed work and discuss its connections to the state-of-the-art.
Summary - proposed approach: The proposed targeted exploration strategy outlined in Algorithm 1 yields a multi-sine exploration input with minimal input energy to generate data from which estimates of uncertain parameters can be derived with a desired error bound (4). The frequencies $\omega_{i}$ of the multi-sine input (6) are predetermined, enabling intuitive tuning based on prior knowledge about the system. The proposed approach assumes energy-bounded disturbances, as commonly considered for data-driven models [fogel1979system, van2020noisy, berberich2022combining]. As the main result, the data-dependent uncertainty bound in Lemma 6 is utilized to derive sufficient conditions in the spectral content of the exploration inputs. The proposed exploration is targeted, as the optimized amplitudes at different frequencies impact both the magnitude and the shape/orientation of the remaining uncertainty after exploration. Furthermore, the proposed exploration strategy is robust, i.e., parametric uncertainty is accounted for by using robust control tools.
Limitations: The proposed exploration strategy relies on the energy-bound of the disturbances $\gamma_{\mathrm{w}}$. The strategy requires solving an SDP iteratively to mitigate suboptimality arising from the convex relaxation procedure. Additionally, the proposed robust exploration strategy is more conservative for large initial uncertainty.
Related works: The derived targeted exploration strategy is similar to, and inspired by [venkatasubramanian2023sequential, umenberger2019robust, ferizbegovic2019learning, barenthin2008identification], however, with a few crucial differences. A targeted exploration method is proposed in [barenthin2008identification] to identify parameters up to a desired accuracy. However, the conditions are not robust to uncertainty and hence iterative experiments are required in practice. In [ferizbegovic2019learning], the proposed targeted exploration strategy assumes independent data, and hence lacks applicability to correlated time-series data from a single trajectory. Furthermore, the methods in [ferizbegovic2019learning] and [umenberger2019robust] do not yield any guarantees for exploration since the uncertainty bounds are approximated in a heuristic way. The method in [venkatasubramanian2023sequential] robustly accounts for parametric uncertainties and provides an a priori guaranteed bound on the uncertainty after exploration. However, all the methods [barenthin2008identification, ferizbegovic2019learning, umenberger2019robust, venkatasubramanian2023sequential] assume i.i.d. Gaussian disturbances. In contrast to these methods, we consider energy-bounded disturbances without assumptions on the distribution or independence of the disturbances. This encompasses a broader class of uncertainties, including those arising from unmodeled dynamics or nonlinearities, and enables the development of a targeted exploration strategy with guarantees. In our proposed targeted exploration method, we quantify and guarantee a priori uncertainty bounds on the parameters obtained from an experiment. Similar to [venkatasubramanian2023sequential], we provide a priori guarantees by robustly accounting for the impact of the uncertain model parameters and disturbances. A notable difference from [venkatasubramanian2023sequential] is the appearance of the term $ZZ^{\top}$ in Theorem 7, which cannot be directly addressed with standard robust control tools and presented significant additional challenges in the present paper.
Application to dual control: The proposed targeted exploration strategy may be utilized to design a robust dual control strategy which guarantees a desired performance for the closed loop after exploration [barenthin2008identification, venkatasubramanian2023sequential]. This can be achieved by co-designing the targeted exploration problem (4.6) with a robust gain-scheduled controller in order to account for the changes in uncertainty during the process of exploration [venkatasubramanian2023sequential]. Such a joint design of a dual controller also allows optimizing closed-loop performance after exploration, ensuring the exploration process reduces only the necessary uncertainty to achieve the desired performance within exploration energy constraints.
5 Numerical Example
In this section, we demonstrate the applicability of the proposed targeted exploration strategy to a nonlinear system using a numerical example. Numerical simulations [^1] The source code for the simulations is available at https:////github.com//jananivenkatasubramanian//NonstochTE were performed on MATLAB using CVX [cvx] in conjunction with the solver MOSEK.
Problem Setup: We consider a chain of two mass-spring-damper systems. The model equations are given by
$m_{1}\ddot{p}_{1}$ $=-(k_{1}+k_{2})p_{1}-(d_{1}+d_{2})\dot{p}_{1}+k_{2}p_{2}+d_{2}\dot{p}_{2}+F_{1}^{\mathrm{nl}},$
$m_{2}\ddot{p}_{2}$ $=k_{2}p_{1}+d_{2}\dot{p}_{1}-k_{2}p_{2}-d_{2}\dot{p}_{2}+F_{2}^{\mathrm{nl}}+u,$ (53)
and
$F_{i}^{\mathrm{nl}}$ $=\beta_{i}\tanh(\alpha_{i}\dot{p}_{i}),$ (54)
with positions $p_{i}$, velocities $\dot{p}_{i}$, masses $m_{i}>0$, spring constants $k_{i}\geq 0$, damping coefficients $d_{i}\geq 0$, nonlinear Coulomb frictions $F_{i}^{\mathrm{nl}}$ with constants $\alpha_{i},\,\beta_{i}\geq 0$, $i\in\{1,2\}$, and input $u\in\mathbb{R}$. We consider $x=[p_{1},\dot{p}_{1},p_{2},\dot{p}_{2}]^{\top}\in\mathbb{R}^{4}$ and an Euler discretization with sampling period $T_{s}=0.5$. The resulting system dynamics correspond to our setup (2) with the disturbances $w_{k}=\left[0,\frac{F_{1}^{\mathrm{nl}}(x_{2})}{m_{1}},0,\frac{F_{1}^{\mathrm{nl}}(x_{4})}{m_{2}}\right]^{\top}$. The disturbances satisfy Assumption 1 with $\gamma_{\mathrm{w}}=T\left(\frac{\beta_{1}^{2}}{m_{1}^{2}}+\frac{\beta_{2}^{2}}{m_{2}^{2}}\right)$. In this example, we consider the following true values for the masses, spring constants and damping coefficients: $m_{1}=1$, $m_{2}=2$, $k_{1}=1$, $k_{2}=1.5$, $d_{1}=0.5$ and $d_{2}=1.1$ (5). We select the coefficients of Coulomb friction as $\alpha_{1}=1$ and $\alpha_{2}=1$ in (54). Later, we vary $\beta_{1}$ and $\beta_{2}$ to study the effect of the disturbance bound $\gamma_{\mathrm{w}}$.
The goal of the proposed targeted exploration strategy is to achieve a desired error bound $D_{\mathrm{des}}^{-1}=10^{0}I_{n_{\phi}}$ on the parameters (4). Furthermore, we select $L=20$ equally-spaced frequencies $\omega_{i}\in\{0,0.05,0.1,0.15,...,0.95\}$. We set $T=100$ and $\epsilon=0.5$. In what follows, we analyse the effectiveness and conservatism of the proposed targeted exploration strategy.
Required input energy $\gamma_{\mathrm{e}}^{2}$ for different energy bounds $\gamma_{\mathrm{w}}$: To study the effect of the energy bound $\gamma_{\mathrm{w}}$ on the required input energy $\gamma_{\mathrm{e}}^{2}$ that ensures the exploration goal, we select the initial uncertainty level as $D_{0}^{-1}=10^{-4}I_{n_{\phi}}$ and the initial estimate as $\hat{\theta}_{0}=\theta_{\mathrm{tr}}+\frac{\|D_{0}\|^{\frac{1}{2}}}{\|\theta_{\mathrm{tr}}\|}\theta_{\mathrm{tr}}$, i.e., on the boundary of $\mathbf{\Theta}_{0}$ (cf. Assumption 2). We run five trials for the following energy bounds $\gamma_{\mathrm{w}}\in\{10^{-4},10^{-3},10^{-2},10^{-1},10^{0}\}$. Each trial comprises: $\mathrm{(i)}$ computing the corresponding constants $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$ using the scenario approach [venkatasubramanian2023sequential, Appendix C] with confidence level $\beta=10^{-10}$, $\mathrm{(ii)}$ executing Algorithm 1 to obtain the exploration inputs (6) and the required input energy $\gamma_{\mathrm{e}}^{2}$.
Figure 1: Illustration of the exploration input energy $\gamma_{\mathrm{e}}^{2}$, in comparison with the disturbance energy bound $\gamma_{\mathrm{w}}$ for the initial uncertainty level $\|D_{0}\|=10^{2}$.
From Figure 1, it can be observed that the input energy $\gamma_{\mathrm{e}}^{2}$ scales roughly linearly with the disturbance energy bound $\gamma_{\mathrm{w}}$. Furthermore, the input energy needs to be significantly larger than the disturbance energy to achieve the desired accuracy of the estimated parameters. In general, $\gamma_{\mathrm{e}}\to 0$ as $\gamma_{\mathrm{w}}\to 0$, i.e., the exploration input energy $\gamma_{\mathrm{e}}$ reduces to zero as the energy of the disturbances $\gamma_{\mathrm{w}}$ reduce to zero.
Conservatism related to the initial uncertainty bound $D_{0}$: To study the effect of the initial uncertainty bound $D_{0}^{-1}$, we select eight uncertainty levels $D_{0}\in\{10^{i}I_{n_{\phi}},5\times 10^{i}I_{n_{\phi}}\}$, i=4,,7i=4,...,7 and $\gamma_{\mathrm{w}}=1$. Based on the energy bound $\gamma_{\mathrm{w}}$, we compute $\beta_{1}=\beta_{2}=\frac{m_{1}m_{2}\sqrt{\gamma_{\mathrm{w}}}}{\sqrt{T(m_{1}^{2}+m_{2}^{2})}}$. Each trial comprises: $\mathrm{(i)}$ computing the initial estimate as $\hat{\theta}_{0}=\theta_{\mathrm{tr}}+\frac{\|D_{0}\|^{\frac{1}{2}}}{\|\theta_{\mathrm{tr}}\|}\theta_{\mathrm{tr}}$, $\mathrm{(ii)}$ computing the corresponding constants $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$ using the scenario approach [venkatasubramanian2023sequential, Appendix C] with confidence level $\beta=10^{-10}$, $\mathrm{(iii)}$ executing Algorithm 1 to obtain the exploration inputs (6), and generating a dataset by applying the exploration input. From the datasets, we compute the estimate of the parameters $\hat{\theta}_{T}$, and $\|G\cdot P\|$, the error bound guaranteed by Lemma 6 with the data-dependent covariance matrix $P$ (12) and scaling $G$ (14).
Figure 2: Illustration of the a posteriori guaranteed bound on the squared error of the parameters $\|G\cdot P\|$, in comparison with the desired bound on the squared error $\|D_{\mathrm{des}}^{-1}\|$ for different initial uncertainty bounds $D_{0}^{-1}$.
From Figure 2, it can be observed that the targeted exploration inputs achieve the exploration goal for all tested initial uncertainty levels $D_{0}$ and energy bound $\gamma_{\mathrm{w}}$, i.e., the guaranteed error bound on the parameters $\|G\cdot P\|$ is lower than the desired error bound $\|D_{\mathrm{des}}^{-1}\|~=~10^{0}$. The guaranteed error bound is significantly lower than the desired error bound due to the inherent conservatism of the proposed strategy which is derived by utilizing worst case bounds on the transfer matrices $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$, and worst case bounds $\bar{W}_{\phi}$, $\bar{W}_{Z}$. This conservatism decreases as the initial uncertainty is reduced, or equivalently, as $D_{0}$ increases. This occurs because smaller initial uncertainty results in tighter and more accurate bounds on $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$, allowing for more efficient and targeted exploration.
Comparison with naive exploration: To highlight the benefits of the proposed method, we compare the targeted exploration method with a naive exploration strategy that uses inputs of equal total energy but with non-optimized amplitudes, i.e., inputs where the energy is uniformly distributed across all frequencies. This baseline does not rely on any model information and illustrates how optimization improves the input design. We select the same eight uncertainty levels $D_{0}\in\{10^{i}I_{n_{\phi}},5\times 10^{i}I_{n_{\phi}}\}$, i=4,,7i=4,...,7 as above, and $\gamma_{\mathrm{w}}=1$. For each trial, we augment the datasets with data generated using the corresponding naive input. From the datasets, we compute the estimate of the parameters $\hat{\theta}_{T}$, and $\|G\cdot P\|$ for the proposed targeted exploration and the baseline. From Figure 3, it can be observed that the proposed targeted exploration method guarantees a lower error bound $\|G\cdot P\|$, by roughly $50\%$, compared to naive exploration. This comparison demonstrates that the proposed targeted input design guarantees lower parameter error under the same energy budget, thereby highlighting the benefits of our approach.
Figure 3: Illustration of the a posteriori guaranteed bound on the squared error of the parameters $\|G\cdot P\|$ for both targeted and naive exploration with same input energy, for different initial uncertainty bounds $D_{0}^{-1}$.
Sensitivity related to the initial estimate $\hat{\theta}_{0}$: To study the sensitivity related to the initial estimate $\hat{\theta}_{0}$, we select three different uncertainty levels $D_{0}\in\{10^{i}I_{n_{\phi}}\}$, $i=4,5,6$. We select $\gamma_{\mathrm{w}}=1$. For each uncertainty level, we run 50 trials. Each trial comprises: $\mathrm{(i)}$ generating a random initial estimate $\hat{\theta}_{0}$ that satisfies Assumption 2, i.e., $(\hat{\theta}_{0}-\theta_{\mathrm{tr}})^{\top}(D_{0}\otimes I_{n_{\mathrm{x}}})(\hat{\theta}_{0}-\theta_{\mathrm{tr}})\leq 1$, $\mathrm{(ii)}$ computing the corresponding constants $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$ using the scenario approach [venkatasubramanian2023sequential, Appendix C] with confidence level $\beta=10^{-10}$, $\mathrm{(iii)}$ executing Algorithm 1 to obtain the exploration inputs (6) and the input energy $\gamma_{\mathrm{e}}^{2}$.
Figure 4: Illustration of the distribution of the input energy $\gamma_{\mathrm{e}}^{2}$ for different initial uncertainty bounds $D_{0}^{-1}$.
From Figure 4, it can be observed that the variability of input energy $\gamma_{\mathrm{e}}^{2}$ is higher for larger initial uncertainty levels $\|D_{0}^{-1}\|$. This variability of the input energy significantly reduces as initial uncertainty levels reduce. This is because small initial uncertainty levels yield a tighter range of bounds on $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$, thereby leading to more consistent exploration input energy.
Sensitivity related to the desired error bound $\|D_{\mathrm{des}}\|$: To study the sensitivity related to the desired uncertainty level $\|D_{\mathrm{des}}\|$, we select six levels of uncertainty $D_{\mathrm{des}}=10^{i}I_{n_{\phi}}$, $i\in\{-4,-3,-2,-1,0,1\}$. We set the initial uncertainty level to $D_{0}=10^{4}I_{n_{\phi}}$, $\gamma_{\mathrm{w}}=1$ and select the initial estimate as $\hat{\theta}_{0}=\theta_{\mathrm{tr}}+\frac{\|D_{0}\|^{\frac{1}{2}}}{\|\theta_{\mathrm{tr}}\|}\theta_{\mathrm{tr}}$. We compute the corresponding constants $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$ using the scenario approach [venkatasubramanian2023sequential, Appendix C] with confidence level $\beta=10^{-10}$. For each desired uncertainty level $\|D_{\mathrm{des}}\|$, we execute Algorithm 1 to obtain the exploration inputs (6) and the input energy $\gamma_{\mathrm{e}}^{2}$.
Figure 5: Illustration of input energy $\gamma_{\mathrm{e}}^{2}$ for different desired uncertainty levels $\|D_{\mathrm{des}}\|$.
From Figure 5, it can be observed that achieving a small uncertainty (large $\|D_{\mathrm{des}}\|$) requires larger input energy. In contrast, in the setting of stochastic disturbances [venkatasubramanian2023sequential], the required input energy scales linearly with $\|D_{\mathrm{des}}\|$. This difference is due to the data-dependent scaling $G$ (14) of the non-falsified set for energy-bounded disturbances (cf. Lemma 6).
Computational time: The average execution time for solving Problem (4.6) over 10 trials is approximately 45 seconds. In general, Problem (4.6) is a standard SDP, and its computational effort scales polynomially with the state dimension $n_{\mathrm{x}}$, control input dimension $n_{\mathrm{u}}$, and the number of frequencies $L$. While this makes moderately-sized problems tractable, the computational effort may become prohibitive for large-scale problems. However, for large-scale problems, problem-specific structures can often be exploited and efficient solvers may be utilized to reduce computational demand [majumdar2020recent]. The simulations were carried out on a system with an AMD Ryzen 7 5700U processor and 16.0 GB RAM.
In summary, the simulation results demonstrate the applicability of the proposed exploration strategy to systems with unmodeled nonlinearities. Given an initial estimate $\hat{\theta}_{0}$ and initial uncertainty bound $D_{0}^{-1}$, the targeted exploration strategy guarantees an a priori error bound on parameters estimated after exploration.
6 Conclusion
In this article, we presented a targeted exploration strategy for linear systems subject to energy-bounded disturbances. We derived LMIs that can robustly guarantee an a priori error-bound on the estimated parameters after exploration. The proposed strategy utilizes multi-sine inputs in selected frequencies and optimized amplitudes to shape the uncertainty bound over the parameters in a targeted manner. We have demonstrated the applicability of the proposed targeted exploration strategy to systems with nonlinearities, while analysing its benefits and inherent conservatism, with a numerical example. Overall, to the best of the authors’ knowledge, this is the first targeted exploration approach that robustly ensures a user-chosen accuracy on the parameters without requiring any independence conditions on the disturbances.
Appendix A Proof of Lemma 9
{pf}
[…]
$\succeq T\begin{bmatrix}D_{\mathrm{des}}^{\frac{1}{2}\top}\left(\left(\sum_{i=1}^{L}\bar{x}(\omega_{i})^{\mathsf{H}}\bar{x}(\omega_{i})\right)\otimes I_{n_{\phi}}\right)D_{\mathrm{des}}^{\frac{1}{2}}&D_{\mathrm{des}}^{\frac{1}{2}\top}\left(\sum_{i=1}^{L}\bar{x}(\omega_{i})^{\top}\left(\bar{\phi}(\omega_{i})^{\mathsf{H}}\otimes I_{n_{\mathrm{x}}}\right)\right)\otimes I_{n_{\phi}}\\ \left(\sum_{i=1}^{L}\bar{x}(\omega_{i})^{\top}\left(\bar{\phi}(\omega_{i})^{\mathsf{H}}\otimes I_{n_{\mathrm{x}}}\right)\right)^{\mathsf{H}}\otimes I_{n_{\phi}}D_{\mathrm{des}}^{\frac{1}{2}}&\left(\left(\sum_{i=1}^{L}\bar{\phi}(\omega_{i})\bar{\phi}(\omega_{i})^{\mathsf{H}}\right)\otimes I_{n_{\mathrm{x}}}\right)\otimes I_{n_{\phi}}\end{bmatrix}$ (55)
The proof is divided into two parts. We first derive (34) and then derive (35).
Part I. From the Parseval-Plancherel identity, we have
[…]
$\succeq T\left(\sum_{i=1}^{L}\bar{\phi}(\omega_{i})\bar{\phi}(\omega_{i})^{\mathsf{H}}\right)$
$=T\left(\bar{\Phi}_{\mathrm{u}}+\bar{\Phi}_{\mathrm{w}}\right)\left(\bar{\Phi}_{\mathrm{u}}+\bar{\Phi}_{\mathrm{w}}\right)^{\mathsf{H}}.$ (56)
By Young’s inequality [caverly2019lmi], for any $\epsilon>0$, we have
$\bar{\Phi}_{\mathrm{u}}\bar{\Phi}_{\mathrm{w}}^{\mathsf{H}}+\bar{\Phi}_{\mathrm{w}}\bar{\Phi}_{\mathrm{u}}^{\mathsf{H}}\succeq-\epsilon\bar{\Phi}_{\mathrm{u}}\bar{\Phi}_{\mathrm{u}}^{\mathsf{H}}-\frac{1}{\epsilon}\bar{\Phi}_{\mathrm{w}}\bar{\Phi}_{\mathrm{w}}^{\mathsf{H}}$ (57)
and hence,
$(\bar{\Phi}_{\mathrm{u}}+\bar{\Phi}_{\mathrm{w}})(\bar{\Phi}_{\mathrm{u}}+\bar{\Phi}_{\mathrm{w}})^{\mathsf{H}}\succeq(1-\epsilon)\bar{\Phi}_{\mathrm{u}}\bar{\Phi}_{\mathrm{u}}^{\mathsf{H}}-\left(\frac{1-\epsilon}{\epsilon}\right)\bar{\Phi}_{\mathrm{w}}\bar{\Phi}_{\mathrm{w}}^{\mathsf{H}}.$ (58)
By inserting Inequality (58) in Inequality (A), we get (34).
Part II. From the Parseval-Plancheral identity, we have (A), which can be written as
$ZZ^{\top}\succeq T\bar{Z}\bar{Z}^{\mathsf{H}}=T\left(\bar{Z}_{\mathrm{u}}+\bar{Z}_{\mathrm{w}}\right)\left(\bar{Z}_{\mathrm{u}}+\bar{Z}_{\mathrm{w}}\right)^{\mathsf{H}}.$ (59)
By Young’s inequality [caverly2019lmi], for any $\epsilon>0$, we have
$\bar{Z}_{\mathrm{u}}\bar{Z}_{\mathrm{w}}^{\mathsf{H}}+\bar{Z}_{\mathrm{w}}\bar{Z}_{\mathrm{u}}^{\mathsf{H}}\succeq-\epsilon\bar{Z}_{\mathrm{u}}\bar{Z}_{\mathrm{u}}^{\mathsf{H}}-\frac{1}{\epsilon}\bar{Z}_{\mathrm{w}}\bar{Z}_{\mathrm{w}}^{\mathsf{H}}$ (60)
and hence,
$(\bar{Z}_{\mathrm{u}}+\bar{Z}_{\mathrm{w}})(\bar{Z}_{\mathrm{u}}+\bar{Z}_{\mathrm{w}})^{\mathsf{H}}\succeq(1-\epsilon)\bar{Z}_{\mathrm{u}}\bar{Z}_{\mathrm{u}}^{\mathsf{H}}-\left(\frac{1-\epsilon}{\epsilon}\right)\bar{Z}_{\mathrm{w}}\bar{Z}_{\mathrm{w}}^{\mathsf{H}}.$ (61)
[…]
Appendix B Proof of Lemma 10
{pf}
The proof is provided in two parts. In the first part, we prove (40), and in the second part, we prove (41).
Part I. In order to determine a bound on $\bar{\Phi}_{\mathrm{w}}$ (31) of the form in (40), we first determine a bound on $W$ (4.2). In particular, we have
$\|W\|=\max_{i=1,\dots,L}\|\bar{w}(\omega_{i})\|$ (62)
and
$\max_{i=1,\dots,L}\|\bar{w}(\omega_{i})\|^{2}\leq\sum_{i=1}^{L}\|\bar{w}(\omega_{i})\|^{2}\leq\sum_{\omega_{i}\in\Omega_{T}}\|\bar{w}(\omega_{i})\|^{2}.$ (63)
From the Parseval-Plancherel identity and the energy-bound on the disturbance (3), we have
[…]
and hence, from (62) and (64), we have
$\|W\|\leq\sqrt{\frac{\gamma_{\mathrm{w}}}{T}}\text{ and }WW^{\mathsf{H}}\preceq\frac{\gamma_{\mathrm{w}}}{T}I_{n_{\mathrm{x}}L}.$ (65)
Starting from (31), we get
$\bar{\Phi}_{\mathrm{w}}\bar{\Phi}_{\mathrm{w}}^{\mathsf{H}}$ $=Y_{\phi,\mathrm{tr}}WW^{\mathsf{H}}Y_{\phi,\mathrm{tr}}^{\mathsf{H}}$
[…]
Part II. From (40), we have
$\|\bar{\Phi}_{\mathrm{w}}\|\leq\sqrt{\frac{\gamma_{\mathrm{w}}}{T}}\|\Gamma_{\phi}\|^{\frac{1}{2}}.$ (66)
Similarly, from (27), we have
$\|\bar{X}_{\mathrm{w}}\|\leq$ $\|Y_{\mathrm{\mathrm{x},tr}}\|\|W\mathbf{1}_{L}\|$ (67)
[…]
Recall that $\bar{Z}_{\mathrm{w}}=\begin{bmatrix}D_{\mathrm{des}}^{\frac{1}{2}\top}\left(\bar{X}_{\mathrm{w}}^{\mathsf{H}}\otimes I_{n_{\phi}}\right)\\ \bar{\Phi}_{\mathrm{w}}\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}$ from (4.2). Note that
$\bar{Z}_{\mathrm{w}}^{\mathsf{H}}\bar{Z}_{\mathrm{w}}=$ $\left(\bar{X}_{\mathrm{w}}\otimes I_{n_{\phi}}\right)D_{\mathrm{des}}^{\frac{1}{2}}D_{\mathrm{des}}^{\frac{1}{2}\top}\left(\bar{X}_{\mathrm{w}}^{\mathsf{H}}\otimes I_{n_{\phi}}\right)$
$+\left(\bar{\Phi}_{\mathrm{w}}^{\mathsf{H}}\bar{\Phi}_{\mathrm{w}}\otimes I_{n_{\mathrm{x}}n_{\phi}}\right).$ (68)
By taking the induced norm of both sides of (B), followed by the application of the triangle inequality, we have
$\|\bar{Z}_{\mathrm{w}}\|^{2}\leq$ $\|D_{\mathrm{des}}^{\frac{1}{2}\top}(\bar{X}_{\mathrm{w}}^{\mathsf{H}}\otimes I_{n_{\mathrm{x}}n_{\phi}})\|^{2}+\|\bar{\Phi}_{\mathrm{w}}\|^{2}$
[…]
Finally, since $\|\bar{Z}_{\mathrm{w}}\bar{Z}_{\mathrm{w}}^{\mathsf{H}}\|=\|\bar{Z}_{\mathrm{w}}^{\mathsf{H}}\bar{Z}_{\mathrm{w}}\|=\|\bar{Z}_{\mathrm{w}}\|^{2}$, from (B) we have
$\bar{Z}_{\mathrm{w}}\bar{Z}_{\mathrm{w}}^{\mathsf{H}}$ $\preceq\|\bar{Z}_{\mathrm{w}}\|^{2}I_{(n_{\phi}+n_{\mathrm{x}}n_{\phi}^{2})}$
$=\left(\frac{\gamma_{\mathrm{w}}}{T}\|\Gamma_{\mathrm{x}}\|\|D_{\mathrm{des}}\|+\frac{\gamma_{\mathrm{w}}}{T}\|\Gamma_{\phi}\|\right)I_{(n_{\phi}+n_{\mathrm{x}}n_{\phi}^{2})}$
[…]
Appendix C Proof of Theorem 14
{pf}
Inequalities (48a)-(48c) imply the exploration goal (4) due to Proposition 13. In what follows, we prove that Inequalities (49a)-(49c) imply Inequalities (48a)-(48c), respectively. In particular, we utilize the matrix S-lemma [boyd2004convex, vanwaarde2022noisy] to account for uncertainties in $V_{\mathrm{x,tr}}$, $V_{\mathrm{\phi,tr}}$, $Y_{\mathrm{x,tr}}$, and $Y_{\mathrm{\phi,tr}}$ satisfying bounds (38), (39). The proof is divided into three parts wherein each part derives (49a), (49b) and (49c), respectively.
Part I. Inequality (48a) can be written as
$\begin{bmatrix}(V_{\mathrm{x,tr}}^{\mathsf{H}}\otimes I_{n_{\phi}})\hat{Z}^{\mathsf{H}}\\ I\end{bmatrix}^{\mathsf{H}}$
$\times\begin{bmatrix}0&\star^{\mathsf{H}}&\\ \underbrace{\begin{bmatrix}(1-\epsilon)D_{\mathrm{des}}^{\frac{1}{2}\top}\left(\mathbf{1}_{L}^{\top}U_{\mathrm{e}}^{\top}\otimes I_{n_{\phi}}\right)\\ 0\end{bmatrix}}_{\star}&-\bar{D}_{1}\end{bmatrix}$
$\times\begin{bmatrix}(V_{\mathrm{x,tr}}^{\mathsf{H}}\otimes I_{n_{\phi}})\hat{Z}^{\mathsf{H}}\\ I\end{bmatrix}$ $\succeq 0.$ (70)
From (36) and (38), we have
$\hat{Z}\left(\left(V_{\mathrm{\mathrm{x},tr}}V_{\mathrm{\mathrm{x},tr}}^{\mathsf{H}}-V_{\mathrm{\mathrm{x},tr}}\hat{V}_{\mathrm{x}}^{\mathsf{H}}-\hat{V}_{\mathrm{x}}V_{\mathrm{\mathrm{x},tr}}^{\mathsf{H}}\right)\otimes I_{n_{\phi}}\right)\hat{Z}^{\mathsf{H}}$
$\preceq$ $\;\hat{Z}\left(\left(\tilde{\Gamma}_{\mathrm{x}}-\hat{V}_{\mathrm{x}}\hat{V}_{\mathrm{x}}^{\mathsf{H}}\right)\otimes I_{n_{\phi}}\right)\hat{Z}^{\mathsf{H}},$ (71)
which can be equivalently written as
$\begin{bmatrix}(V_{\mathrm{x,tr}}^{\mathsf{H}}\otimes I_{n_{\phi}})\hat{Z}^{\mathsf{H}}\\ I\end{bmatrix}^{\mathsf{H}}$
$\times\begin{bmatrix}-I&(\hat{V}_{\mathrm{x}}^{\mathsf{H}}\otimes I_{n_{\phi}})\hat{Z}^{\mathsf{H}}\\ \hat{Z}(\hat{V}_{\mathrm{x}}\otimes I_{n_{\phi}})&\hat{Z}((\tilde{\Gamma}_{\mathrm{x}}-\hat{V}_{\mathrm{x}}\hat{V}_{\mathrm{x}}^{\mathsf{H}})\otimes I_{n_{\phi}})\hat{Z}^{\mathsf{H}}\end{bmatrix}$ (72)
$\times\begin{bmatrix}(V_{\mathrm{x,tr}}^{\mathsf{H}}\otimes I_{n_{\phi}})\hat{Z}^{\mathsf{H}}\\ I\end{bmatrix}$ $\succeq 0.$
By using the matrix S-lemma [boyd2004convex, vanwaarde2022noisy], Inequality (C) holds for all $V_{\mathrm{x,tr}}$ satisfying Inequality (C), if $S_{\textnormal{exp-1}}(\epsilon,\tau_{1},U_{\mathrm{e}},\hat{Z},\hat{V}_{\mathrm{x}},\tilde{\Gamma}_{\mathrm{x}},\tilde{D}_{\mathrm{des}},\bar{D}_{1})~\succeq~0$ (49a) holds with $\tau_{1}\geq 0$.
Part II. Inequality (48b) can be written as
$\begin{bmatrix}\begin{bmatrix}0\\ V_{\mathrm{\phi,tr}}\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}^{\mathsf{H}}\\ I\end{bmatrix}^{\mathsf{H}}$
$\times\begin{bmatrix}0&\underbrace{(1-\epsilon)(U_{\mathrm{e}}\otimes I_{n_{\mathrm{x}}n_{\phi}})\hat{Z}^{\mathsf{H}}}_{\star}\\ \star^{\mathsf{H}}&-(1-\epsilon)\hat{Z}\hat{Z}^{\mathsf{H}}-\left(\tfrac{1-\epsilon}{\epsilon}\right)\bar{W}_{Z}-\bar{D}_{2}\end{bmatrix}$
$\times\begin{bmatrix}\begin{bmatrix}0\\ V_{\mathrm{\phi,tr}}\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}^{\mathsf{H}}\\ I\end{bmatrix}$ $\succeq 0.$ (73)
From (36) and (38), we have
$(V_{\mathrm{\phi,tr}}V_{\mathrm{\phi,tr}}^{\mathsf{H}}-V_{\mathrm{\phi,tr}}\hat{V}_{\phi}^{\mathsf{H}}-\hat{V}_{\phi}V_{\mathrm{\phi,tr}}^{\mathsf{H}}+\hat{V}_{\phi}\hat{V}_{\phi}^{\mathsf{H}})\otimes I_{n_{\mathrm{x}}n_{\phi}}$
$\preceq$ $\;\tilde{\Gamma}_{\phi}\otimes I_{n_{\mathrm{x}}n_{\phi}},$ (74)
which can be equivalently written as
$\begin{bmatrix}\begin{bmatrix}0\\ V_{\mathrm{\phi,tr}}\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}^{\mathsf{H}}\\ I\end{bmatrix}^{\mathsf{H}}$
$\times\begin{bmatrix}-I&\begin{bmatrix}0\\ \hat{V}_{\mathrm{\phi}}\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}^{\mathsf{H}}\\ \begin{bmatrix}0\\ \hat{V}_{\mathrm{\phi}}\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}&\begin{bmatrix}0&0\\ 0&(\tilde{\Gamma}_{\phi}-\hat{V}_{\mathrm{\phi}}\hat{V}_{\mathrm{\phi}}^{\mathsf{H}})\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}\end{bmatrix}$ (75)
$\times\begin{bmatrix}\begin{bmatrix}0\\ V_{\mathrm{\phi,tr}}\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}^{\mathsf{H}}\\ I\end{bmatrix}$ $\succeq 0.$
By using the matrix S-lemma, Inequality (C) holds for all $V_{\mathrm{\phi,tr}}$ satisfying Inequality (C), if $S_{\textnormal{exp-2}}(\epsilon,\tau_{2},U_{\mathrm{e}},\hat{Z},\bar{W}_{Z},\hat{V}_{\phi},\tilde{\Gamma}_{\phi},\tilde{D}_{\mathrm{des}},\bar{D}_{2})\succeq 0$ (49b) holds with $\tau_{2}\geq 0$.
Part III. Inequality (48c) can be written as
$\begin{bmatrix}\begin{bmatrix}V_{\phi,\mathrm{tr}}\\ 0\end{bmatrix}^{\mathsf{H}}\\ I\end{bmatrix}^{\mathsf{H}}$
$\times\begin{bmatrix}(1-\epsilon)U_{\mathrm{e}}U_{\mathrm{e}}^{T}&0\\ 0&\begin{bmatrix}-\left(\tfrac{1-\epsilon}{\epsilon}\right)\bar{W}_{\phi}-\frac{\gamma_{\mathrm{w}}}{T}D_{\mathrm{des}}&0\\ 0&0\end{bmatrix}-\bar{D}_{3}\end{bmatrix}$
$\times\begin{bmatrix}\begin{bmatrix}V_{\phi,\mathrm{tr}}\\ 0\end{bmatrix}^{\mathsf{H}}\\ I\end{bmatrix}$
$\succeq$  $\;0.$ (76)
From (36) and (38), we have
$V_{\mathrm{\phi,tr}}V_{\mathrm{\phi,tr}}^{\mathsf{H}}-V_{\mathrm{\phi,tr}}\hat{V}_{\phi}^{\mathsf{H}}-\hat{V}_{\phi}V_{\mathrm{\phi,tr}}^{\mathsf{H}}+\hat{V}_{\phi}\hat{V}_{\phi}^{\mathsf{H}}\preceq\tilde{\Gamma}_{\phi},$ (77)
which can be equivalently written as
$\begin{bmatrix}\begin{bmatrix}V_{\mathrm{\phi,tr}}\\ 0\end{bmatrix}^{\mathsf{H}}\\ I\end{bmatrix}^{\mathsf{H}}\begin{bmatrix}-I&\begin{bmatrix}\hat{V}_{\mathrm{\phi}}\\ 0\end{bmatrix}^{\mathsf{H}}\\ \begin{bmatrix}\hat{V}_{\mathrm{\phi}}\\ 0\end{bmatrix}&\begin{bmatrix}(\tilde{\Gamma}_{\phi}-\hat{V}_{\mathrm{\phi}}\hat{V}_{\mathrm{\phi}}^{\mathsf{H}})&0\\ 0&0\end{bmatrix}\end{bmatrix}$
$\times\begin{bmatrix}\begin{bmatrix}V_{\mathrm{\phi,tr}}\\ 0\end{bmatrix}^{\mathsf{H}}\\ I\end{bmatrix}$ $\succeq 0.$ (78)
By using the matrix S-lemma, Inequality (C) holds for all $V_{\mathrm{\phi,tr}}$ satisfying Inequality (C), if the following holds with $\tau_{3}\geq 0$:
$\begin{bmatrix}(1-\epsilon)U_{\mathrm{e}}U_{\mathrm{e}}^{T}&0\\ 0&\begin{bmatrix}-\left(\tfrac{1-\epsilon}{\epsilon}\right)\bar{W}_{\phi}-\frac{\gamma_{\mathrm{w}}}{T}D_{\mathrm{des}}&0\\ 0&0\end{bmatrix}-\bar{D}_{3}\end{bmatrix}$
$-\tau_{3}\begin{bmatrix}-I&\begin{bmatrix}\hat{V}_{\mathrm{\phi}}\\ 0\end{bmatrix}^{\mathsf{H}}\\ \ \begin{bmatrix}\hat{V}_{\mathrm{\phi}}\\ 0\end{bmatrix}&\begin{bmatrix}(\tilde{\Gamma}_{\phi}-\hat{V}_{\mathrm{\phi}}\hat{V}_{\mathrm{\phi}}^{\mathsf{H}})&0\\ 0&0\end{bmatrix}\end{bmatrix}$ $\succeq 0.$ (79)
From Lemma 12, we have
$U_{\mathrm{e}}U_{\mathrm{e}}^{T}\succeq U_{\mathrm{e}}\hat{U}^{\top}+\hat{U}U_{\mathrm{e}}^{\top}-\hat{U}\hat{U}^{\top}.$ (80)
Inserting Inequality (80) in Inequality (C) yields $S_{\textnormal{exp-3}}(\epsilon,\tau_{3},U_{\mathrm{e}},\hat{U},\hat{W}_{\phi},\hat{V}_{\phi},\tilde{\Gamma}_{\phi},D_{\mathrm{des}},\bar{D}_{3})\succeq 0$ (49c).
[…]